OpenStudy (anonymous):

solve equation x^2e^2x+2xe^2x=8e^2x please help!

OpenStudy (anonymous):

\[x ^{2}e ^{2x}+2xe ^{2x}=8e ^{2x}\]

OpenStudy (muscrat123):

Ah, for this one: x^2e^2x + 2xe^2x = 8e^2x Just divide everything by e^2x: x^2 + 2x = 8 x^2 + 2x - 8 = 0 Looks better, doesn't it? Foil! (x-2)(x+4)=0 x=2 or x=-4

OpenStudy (anonymous):

Thank you!

OpenStudy (muscrat123):

np